OpenStudy (anonymous):

integral of sec4xtan4xdx

OpenStudy (anonymous):

(1/4)sec(4x)

myininaya (myininaya):

let u=4x so du=4 dx or you can say du/4=dx sp we have 1/4 * int(secutanu,u) =1/4 * secu+c (since the derivative of secu= secutanu) =1/4 * sec(4x)+c

OpenStudy (anonymous):

*sigh* forgot the C

myininaya (myininaya):

how could you!!!!

myininaya (myininaya):

most important part lol